The time period of a satellite of the earth is 10 hours. If the separation between the earth and the satellite is increased to 4 times the previous value, then what will be the new time period of the satellite?
 
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From Kepler’s third law of planetery motion
T^2  =  k* R^3
 
Therfore,
(10)^2 = k.R^3.....................(1)
after four time distance,
T^2 = k.(4R)^3
T^2 = k.64R^3 …..............(2)
Dividing eqns.  (1)  and  (2)
(10/T)^2 = (1/64)
10/T = 1/8
T = 80hours.